Potential difference between two conductors?

A coaxial cable consists of a long cylindrical conductor of radius a=1 mm
concentric with a thin cylindrical shell of larger radius b=5 mm. If the central conductor has a charge per unit length ?=Q/L=1 micro(C)/m uniformly distributed on itssurface, what is the potential difference between the inner and outer conductors?
Answer #1
GivenRadius of inner cylinder a = 1 mm= ( 1 mm ) ( 10-3 m / 1 mm )= 10-3 mRadius of outer cylinder b = 5 mm=(5 mm ) ( 10-3 m / 1 mm )= 5*10-3 mCharge distributed on inner cylinderλ = Q / L= 1 μC / m= ( 1 μC / m ) ( 10-6 C / 1 μC )= 10-6 C / mElectric potential difference between coaxial cylinderV = ( Q / 2πεo L )( ln ( b / a )= ( λ / 2πεo ) ln ( b / a)=[ ( 10-6 C / m ) / ( 2π*8.85*10-12 C 2 / N. m2 ) ] ln ( 5 mm / 1 mm )=28943 V
answered by: ajones